In a geometric progression the sum of 2nd and 4th terms is 30. The difference of
6th and 2nd terms is 90. Find the 8th term of a geometric progression, whose
common ratio is greater than 1.

Answer by richwmiller(17219)

Tn = t * r^(n - 1)
30 = t * r^(1)+t * r^(3),
90 = t * r^(5)-t * r^(1)
r = 2, t = 3
Tn = t * r^(n - 1)
T8 = 3 * 2^(7)
T8 = 384